### v2

# HandSSCA: 3D Hand Mesh Reconstruction with State Space Channel Attention from RGB images

Abstract
Reconstructing a hand mesh from a single RGB image is a challenging task because hands are often occluded by objects. Most previous works attempted to introduce more additional information and adopt attention mechanisms to improve 3D reconstruction results, but it would increased computational complexity. This observation prompts us to propose a new and concise architecture while improving computational efficiency. In this work, we propose a simple and effective 3D hand mesh reconstruction network HandSSCA, which is the first to incorporate state space modeling into the field of hand pose estimation. In the network, we have designed a novel state space channel attention module that extends the effective sensory field, extracts hand features in the spatial dimension, and enhances hand regional features in the channel dimension. This design helps to reconstruct a complete and detailed hand mesh. Extensive experiments conducted on well-known datasets featuring challenging hand-object occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandSSCA achieves state-of-the-art performance while maintaining a minimal parameter count.
1 Introduction
Reconstruction of 3D hand pose from a single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction. In order to reduce the cost of use in above areas, 3D hand pose estimation based on RGB images has attracted much attention due to its low cost and computational friendliness, which provides new solutions for practical applications. Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field. These works focus on issues related to hand-object occlusion and the loss of local information. Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when there is a lack of prior knowledge.
The previous methods can be broadly categorized into two groups, one is by incorporating additional a priori knowledge. For example, Zimm et al. zimmermann2017learning obtained hand region features via a hand segmentation network. And Wang et al. Wang_2023 added keypoint coordinates as strong a priori information to constrain hand pose. These methods can enhance pose estimation performance, but they lack essential a priori information in practical applications. Another category is to learn the intrinsic connection between image features through deep networks, Park et al. park2022handoccnet extracts the complete hand information by distinguishing the features. jiao et al. https://doi.org/10.1049/ell2.13040 uses the Transformer architecture to obtain the global information and expects to learn the hand feature mapping from the occluded part. These methods tend to mine the association between image features, but there are still some limitations. First, the Attention mechanism leads to excessive computational complexity. Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure.1(b).
Based on the above observations, we propose a simple and effective 3D hand reconstruction network HandSSCA. This framework for the first time incorporates state space model liu2024vmamba; ma2024umamba; ruan2024vmune into the field of hand pose estimation. At the same time, combining the idea of spatial channel feature fusion hu2019squeezeandexcitation, State Space Channel Attention (SSCA) was constructed, which emphasizes the contribution of channel dimension features to image features, as shown in Fig.3. This not only increases the range of effective receptive fields in the network, but also effectively avoids the computational complexity of traditional attention square steps.Specifically, our main contributions are as follows:
(1) We propose a hand pose estimation model called HandSSCA, which introduces the concept of state space model to the field of hand pose estimation for the first time. This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
(2) We propose a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning. And the state space channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
(3) Our method achieves state-of-the-art on three datasets, FREIHAND, DEXYCB and HO3D, with a small number of parameters.
Figure 1: Effective receptive fields (from left to right, FPN, Self-Attention, State Space Attention(SSA) and State Space Channel Attention(SSCA), respectively).
2 Related work
2.1 State Space Model
With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences. In the field of natural language processing, Albert Gu et al. gu2022efficiently proposed the S4 model to optimize the state space model (SSM) by simulation in order to solve the problems of computational efficiency and long sequence dependency. However, the S4 model would process all inputs in exactly the same way, which is not in line with the practical intuition that each sequence should receive different attention. Therefore, Albert Gu et al. gu2023mamba introduced a selection mechanism to propose the Mamba architecture (S6) to learn the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences. Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to linear. Influenced by this idea, Yue Liu et al. ma2024umamba introduced it to vision by combining the VIT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity. This approach achieves linear complexity without sacrificing the global receptive field, as shown in Figure.1(c). In addition, some researchers ruan2024vmunet; hu2019squeezeandexcitation have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information.
2.2 Additional information based Methods
It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations. In recent years, some scholars Wang_2023; jiang2023a2jtransformer combined human a priori knowledge to guide the network to focus on the target region, Wang et al. Wang_2023 proposed HandGCAT, which learns the a priori knowledge of the hand joint positions through the Knowledge-Guided Graph Convolution module, and then through the Cross-Attention Transformer to guide the network to inject joint information into the occluded region. In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose, therefore, Qi et al. qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
2.3 Attention based Methods
Attentional mechanisms have achieved great success in the field of deep learning, thanks to their ability to map long-term contextual relevance, and in the task of hand pose estimation, occluded hand features can be learned through global mapping, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing. In establishing the global context of hand-object interaction, Pavlakos et al. pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the VIT architecture.Cho et al. cho2022crossattention obtain the features related to the region of interest from the occluded portion through cross-attention. Park et al. park2022handoccnet make full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
Figure 2: Overall architecture of HandSSCA including Trunk, SSCA and Regressor.Our HandSSCA uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer. It then uses SSCA to obtain hand region features by pairwise spatial and channel scanning. Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters. The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
3 Method
3.1 Overview
The overall network architecture is shown in Figure.2 and consists of Backbone, SSCA and Regressor components. Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$. First, the image information is extracted by ResNet-50 and the multi-scale features are fused layer by layer to get the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$. Then, it is fed into the SSCA module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain hand features $H\in\mathbb{R}^{256\times 32\times 32}$. Finally, the hand features $H$ are passed to Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates and output the MANO parameters through the FC layer to obtain the final 3D hand mesh.
3.2 Backbone
In the feature extraction part we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature. Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$. In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling. Then, up-sampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of features of size $64\times 64\times 256$. Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the image information through average pooling.
3.3 State Space Channel Attention
State Space Models.
S4ND nguyen2022s4nd was the first study to introduce state space models into the field of vision by enabling it to receive 2D image information through a simple extension of the S4 model. The S6 module proposed by Gu et al. gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selectable scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity. The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{C}^{N\times N},B,C\in\mathbb{C}^{N},D\in\mathbb{C}^{1}$ with the following equation:
$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\\ \mathrm{y(t)=Ch(t)+Dx(t)}$ (1)
In order to achieve computationally efficient operations zimmermann2019freihand, the above ordinary differential equations need to be transformed into discrete functions. Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times D}}$, the discretization is performed following the following rule:
$h_{i}=\overline{A}h_{i-1}+\overline{B}x_{i},\\ y_{i}=\overline{C}h_{i}+\overline{D}x_{i},\\ \overline{A}=e^{\Delta A},\\ \overline{B}=(e^{\Delta A}-I)A^{-1}B,\\ \overline{C}=C\$ (2)
where $\mathrm{B,C\in\mathbb{C}^{D\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{D}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series ref16:
$B=(e^{\Delta A}-I)A^{-1}B\approx(\Delta A)(\Delta A)^{-1}\Delta B=\Delta B$ (3)
Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B\in\mathbb{R}^{B\times L\times N}\quad,C\in\mathbb{R}^{B\times L\times N}$, and $\Delta\in\mathbb{R}^{B\times L\times D}$ can be obtained from inputs of size $X\in\mathbb{R}^{B\times L\times D}$. this means that S6 can sense contextual information from the inputs and update the weights.
Space select.
Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module. The 1-D convolution is extended to 2-D convolution in S4ND gu2021combining, and this approach leads to the loss of the ability to dynamically update weights based on context. In VMamba liu2024vmamba, rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module. This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane. Then, the sequences from each of the four directions are passed through the S6 block to compute the feature map. Finally, the four sequences are merged to obtain a new sequence of the same size as the input.
Channel select.
Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba. However, in deep image features, channels contain richer and diverse contextual information, which is mostly ignored by these methods. Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Figure.3. We perform scan expansion and scan fusion on the $H\times C$ plane. First, the input features are expanded on the $H\times C$ plane and scanned from four directions to obtain the sequence $C_{1},C_{2},C_{3},C_{4}$. Then, they are fed into the S6 block to compute the contextual correlation in the sequence by Eq. (3), respectively. Finally, the sequences are merged and reconstructed into image feature maps of the same size.
Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information. In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field. We visualized the range of effective receptive fields for different attentions, as shown in Figure.1. As can be seen in Figure.1(d), this novel state space channel attention is able to maintain a larger and deeper receptive field range and global feature mapping capability with a reduced number of parameters and linear complexity. This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
Figure 3: SSCA module framework, we send the image features to parallel spatial and channel branches for scanning, and then process the spatial and channel resultant features and merge them through S6 block respectively, and finally sum them to get the 2D feature map.
3.4 Regressor
In Regressor, we input the hand attention features obtained from SSMA. First, a 2D heat map of each joint $H$ is output through a simple Conv cascade. Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks [34]. Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO. At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
To train our proposed HandSSCA hand pose estimation model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
$\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|+\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$ (4)
Where $*^{\prime}$ denotes the true value, * denotes the predicted value, and $\alpha_{\theta}=10,\alpha_{\beta}=0.1,\alpha_{2d}=1\mathrm{e}2,\alpha_{3d}=1\mathrm{e}4,\alpha_{V}=1\mathrm{e}4$. the optimal solution is obtained by minimizing this objective function for training.
Table 1: COMPARISON WITH STATE-OF-THE-ART METHODS ON FREIHAND.
Method Params↓ PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Hasson et al.Hasson2019LearningJR 33.5M 13.3 13.3 0.429 0.907
ExPoseChoutas2020MonocularEB 11.8 12.2 0.484 0.918
Kulon et al.kulon2020weaklysupervised 519.0M 8.4 8.6 0.614 0.966
I2L-MeshNetmoon2020i2lmeshnet 136.8M 7.4 7.6 0.681 0.973
Pose2Meshchoi2021pose2mesh 75.0M 7.4 7.6 0.683 0.973
Tang et al.9710830 156.5M 6.7 6.7 0.724 0.981
Junhyeong et al.cho2022crossattention 153.0M 6.5 0.982
HandSSCA(Ours) 31.8M 7.6 7.0 0.690 0.982
Table 2: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
Method Params↓ MPJPE↓ PA-MPJPE↓
Spurr et al.spurr2020weakly 17.34 6.83
METROlin2021endtoend 229.5M 15.24 6.99
Liu et al.liu2021semisupervised 34.9M 15.28 6.58
HandOccNetpark2022handoccnet 39.3M 14.04 5.80
HandGCATWang_2023 75.8M 13.76 5.60
HandSSCA(Ours) 31.8M 12.81 5.52
Table 3: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
Method Params↓ PA-MPJPE↓ PA-MPJPE AUC↑ PA-MPVPE↓ PA-MPVPE AUC↑ F@5↑ F@15↑
I2L-MeshNetmoon2020i2lmeshnet 136.8M 11.2 0.775 13.9 0.722 0.409 0.932
Hasson et al.hasson2020leveraging 13.2M 11.0 0.780 11.2 0.777 0.464 0.939
Hampali et al.hampali2020honnotate 10.7 0.788 10.6 0.790 0.506 0.942
METROlin2021endtoend 229.5M 10.4 0.792 11.1 0.779 0.484 0.946
Liu et al.liu2021semisupervised 34.9M 9.9 0.803 9.5 0.810 0.528 0.956
I2UV-HandNetchen2021i2uvhandnet 9.9 0.804 10.1 0.799 0.500 0.943
ArtiBoostli2022artiboost 25.3M 11.4 0.773 10.9 0.782 0.488 0.944
KeypointHampali2021KeypointTS 56.5M 10.8 0.786
HandSSCA(Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.955
4 Experiments
4.1 Implementation details
All implementations were done in PyTorch. We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 learning rate to train with a batch size of 32. Both our training and testing were performed on a server with a CPU15 vCPU AMD and an RTX A5000 (24GB). All other details will be provided in our code.
4.2 Datasets and evaluation metrics
FREIHAND.
This dataset zimmermann2019freihand presents the first massively diverse hand dataset containing 32,560 training samples with the possibility of replacing four backgrounds. Also, there are 3960 test samples under occlusion and poor lighting.
DEXYCB.
This dataset chao2021dexycb is a large-scale dataset for object grasping and contains 582K image frames grasped on 20 YCB objects. We use the official “S0” segmentation to evaluate the right hand pose.
HO3D.
This dataset hampali2020honnotate is a hand-object interaction dataset with severe occlusion and contains 66,034 training samples and 11,524 test samples. The dataset provides RGB images with MANO-based hand joints and meshes and camera parameters. The test set provides only the annotations of the RGB images and the hand bounding box, and the test results can be evaluated through an online submission system.
Evaluation metrics.
For FreiHand, we report PA-MPJPE and PA-MPVPE in mm as well as F-scores. For HO3D we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server. For DexYCB we report MPJPE and PA-MPJPE in mm.
Table 4: ABLATION RESULTS FOR THE FREIHAND.
Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Baseline 8.0 7.4 0.669 0.977
Baseline+Self-Attention 7.7 7.0 0.688 0.981
Baseline+SSCA 7.6 7.7 0.690 0.982
Table 5: ABLATION RESULTS FOR THE DEXYCB.
Method MPJPE↓ PA-MPJPE↓
Baseline 14.07 5.78
Baseline+Self-Attention 14.40 5.81
Baseline+SSCA 12.81 5.52
Table 6: ABLATION RESULTS FOR THE HO3D.
Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Baseline 9.8 9.7 0.516 0.954
Baseline+Self-Attention(4.6M) 10.1 10.1 0.502 0.947
Baseline+SSA(0.57M) 9.8 9.8 0.521 0.952
Baseline+SSCA(0.58M) 9.5 9.5 0.518 0.955
4.3 Comparisons with the State-of-the-art Methods
We compare our proposed method with several recent state-of-the-art methods on three datasets, FreiHand, DEXYCB, and HO3D, as shown in Tables 1-3. As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830. Moreover, its number of parameters shrinks by a factor of nearly 5 compared to models with higher accuracy cho2022crossattention. On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods Hampali2021KeypointTS. Meanwhile, when the number of parameters is similar to the method of Liu et al. liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced. To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023. In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included. Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Figure.4.
4.4 Ablation Study
In order to validate the effectiveness of the HandSSCA method, we further analyzed the different attention modules. The results of the ablation experiments on different datasets are shown in Tables 4-6. On the FreiHand and DEXYCB datasets, we designed three experiments, Baseline, Adding Self-Attention Module, and Adding SSCA Module, as a way to validate that our proposed SSCA is able to expand the effective receptive field while enhancing the hand features. On the HO3D dataset, we designed four experiments, which are Baseline, adding the self-attention module, adding the traditional state space attention module (SSA) and adding the state space channel attention module (SSCA). In this way, we demonstrate that the state space model with the addition of channel information can enhance the ability to extract hand features in the occluded state while having a wider effective receptive field.
From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information. In Figure.5, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges. The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent. The SSCA module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes. Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
In Table 6, it can be found that after the introduction of the state selection of the channel plane, the number of parameters of the model is only slightly increased by 0.01M, however, a significant improvement of nearly 2% is realized in the two performance metrics of PA-MPJPE and PA-MPVPE. This is due to the fact that the state selection in the channel plane enhances the selection and fusion of image features, and is able to extract more important channel information. The current state-space model focuses more on multi-directional scanning in the spatial plane, which is the reason for the 0.3% improvement in the F@5 metrics when using SSA. Using only the spatial plane for state selection, each pixel will focus more on neighboring pixels. However, observing other metrics will show that the channel dimension in the image will then contain richer feature information, which happens to be ignored by SSA.
Figure 4: Qualitative results for the three datasets. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by SSCA, the heat map of keypoints in Regressor, and the hand mesh.
Figure 5: Attention characterization map. It is worth noting that the SSCA module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region. And the method expands the range of the receptive field so that it can capture the global feature mapping. (Orange is the occluded region and red is the hand interaction edge region)
5 Conclusion
In order to effectively mitigate the challenges of the hand pose estimation task under occlusion conditions, we introduce state-space modeling into the field of hand pose estimation for the first time and propose the HandSSCA method. The method expands the effective receptive field and acquires hand features under occlusion by parallel cross-scanning in the spatial plane and the channel plane. Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection. This novel attention architecture is able to acquire sufficient global information and expand the effective sensory field while keeping the number of parameters small. Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations. Meanwhile, we expect this spatial and channel-parallel state-selective attention architecture to make progress in different downstream tasks.

### v3

@@
-HandSSCA: 3D Hand Mesh Reconstruction with State Space Channel Attention from RGB images
+HandS3C: 3D Hand Mesh Reconstruction with State Space Spatial Channel Attention from RGB images
 Abstract
-Reconstructing a hand mesh from a single RGB image is a challenging task because hands are often occluded by objects.
-Most previous works attempted to introduce more additional information and adopt attention mechanisms to improve 3D reconstruction results, but it would increased computational complexity.
-This observation prompts us to propose a new and concise architecture while improving computational efficiency.
-In this work, we propose a simple and effective 3D hand mesh reconstruction network HandSSCA, which is the first to incorporate state space modeling into the field of hand pose estimation.
-In the network, we have designed a novel state space channel attention module that extends the effective sensory field, extracts hand features in the spatial dimension, and enhances hand regional features in the channel dimension.
-This design helps to reconstruct a complete and detailed hand mesh.
-Extensive experiments conducted on well-known datasets featuring challenging hand-object occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandSSCA achieves state-of-the-art performance while maintaining a minimal parameter count.
+Reconstructing the hand mesh from one single RGB image is a challenging task because hands are often occluded by other objects.
+Most previous works attempt to explore more additional information and adopt attention mechanisms for improving 3D reconstruction performance, while it would increase computational complexity simultaneously.
+To achieve a performance-reserving architecture with high computational efficiency, in this work, we propose a simple but effective 3D hand mesh reconstruction network (i.e., HandS3C), which is the first time to incorporate state space model into the task of hand mesh reconstruction.
+In the network, we design a novel state-space spatial-channel attention module that extends the effective receptive field, extracts hand features in the spatial dimension, and enhances regional features of hands in the channel dimension.
+This helps to reconstruct a complete and detailed hand mesh.
+Extensive experiments conducted on well-known datasets facing heavy occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandS3C achieves state-of-the-art performance while maintaining a minimal parameters.
 1 Introduction
-Reconstruction of 3D hand pose from a single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction.
-In order to reduce the cost of use in above areas, 3D hand pose estimation based on RGB images has attracted much attention due to its low cost and computational friendliness, which provides new solutions for practical applications.
+The reconstruction of 3D hand mesh from one single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction.
+3D hand mesh reconstruction based on RGB images for the reconstruction has attracted much attention due to its low cost and computational friendliness.
 Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field.
-These works focus on issues related to hand-object occlusion and the loss of local information.
-Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when there is a lack of prior knowledge.
-The previous methods can be broadly categorized into two groups, one is by incorporating additional a priori knowledge.
-For example, Zimm et al.
+These works focus on important issues related to hand-object occlusion and the loss of local information.
+Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when the prior knowledge is not sufficient.
+The previous methods can be broadly categorized into two categories.
+One is to incorporate additional prior knowledge.
+To cite few, Zimm et al.
 zimmermann2017learning obtained hand region features via a hand segmentation network.
-And Wang et al.
-Wang_2023 added keypoint coordinates as strong a priori information to constrain hand pose.
-These methods can enhance pose estimation performance, but they lack essential a priori information in practical applications.
+Wang et al.
+Wang_2023 added keypoint coordinates as strong prior information to constrain hand pose.
+These methods can enhance pose reconstruction performance, but lack essential prior information in practical applications.
 Another category is to learn the intrinsic connection between image features through deep networks, Park et al.
-park2022handoccnet extracts the complete hand information by distinguishing the features.
+park2022handoccnet extracted the complete hand information by distinguishing the features.
 Jiao et al.
-https://doi.org/10.1049/ell2.13040 uses the Transformer architecture to obtain the global information and expects to learn the hand feature mapping from the occluded part.
+https://doi.org/10.1049/ell2.13040 used the Transformer architecture to obtain the global information and expect to learn the hand feature mapping from the occluded part.
 These methods tend to mine the association between image features, but there are still some limitations.
 First, the attention mechanism leads to excessive computational complexity.
-Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure.1(b).
-Based on the above observations, we propose a simple and effective 3D hand reconstruction network HandSSCA.
-This framework for the first time incorporates state space model liu2024vmamba; ma2024umamba; ruan2024vmune into the field of hand pose estimation.
-At the same time, combining the idea of spatial channel feature fusion hu2019squeezeandexcitation, State Space Channel Attention (SSCA) was constructed, which emphasizes the contribution of channel dimension features to image features, as shown in Fig.3.
-This not only increases the range of effective receptive fields in the network, but also effectively avoids the computational complexity of traditional attention square steps.Specifically, our main contributions are as follows:
-(1) We propose a hand pose estimation model called HandSSCA, which introduces the concept of state space model to the field of hand pose estimation for the first time.
+Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure.
+1 (b).
+Based on the above observations, we propose a simple but effective 3D hand mesh reconstruction network called as HandS3C.
+This framework incorporates the state space model liu2024vmamba; ma2024umamba; ruan2024vmunet into the field of hand mesh reconstruction for the first time.
+At the same time, combining the idea of spatial-channel feature fusion hu2019squeezeandexcitation, State Space Spatial Channel (S3C) attention module is constructed to emphasize the contribution of channel-dimensional features.
+This not only increases the range of effective receptive fields in the network, but also effectively avoids the square-step computational complexity of traditional attention.
+Specifically, our main contributions are summarized as follows:
+(1) We design a 3D hand mesh reconstruction network called as HandS3C, which introduces the state space model to the task of hand mesh reconstruction for the first time.
 This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
-(2) We propose a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning.
-And the state space channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
+(2) We propose a spatial and channel-wise parallel scanning approach to compensate for the missing channel information in planar scanning.
+The state-space spatial-channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
 (3) Our method achieves state-of-the-art on three datasets, FREIHAND, DEXYCB and HO3D, with a small number of parameters.
-Figure 1: Effective receptive fields (from left to right, FPN, Self-Attention, State Space Attention(SSA) and State Space Channel Attention(SSCA), respectively).
+Figure 1: Effective receptive fields (from left to right: FPN, Self-Attention, State Space Attention (SSA) and State-Space Spatial-Channel Attention (S3C)).
 2 Related work
+2.1 Additional Information based Methods
+It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations.
+In recent years, some scholars Wang_2023; jiang2023a2jtransformer incorporate human prior knowledge to guide the network to focus on the target region, Wang et al.
+Wang_2023 proposed HandGCAT, which learns the prior knowledge of the hand joint positions through the knowledge-guided graph convolution module, and then through the cross-attention Transformer to guide the network to inject joint information into the occluded region.
+In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose.
+To end this, Qi et al.
+qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
+2.2 Attention based Methods
+Attention mechanisms have achieved great success in the field of deep learning, due to their ability to map long-term contextual relevance.
+In the task of hand mesh reconstruction, occluded hand features can be learned through global mapping with various attention modules, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing.
+In establishing the global context of hand-object interaction, Pavlakos et al.
+pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the Vision Transformer (ViT) architecture.
+Cho et al.
+cho2022crossattention obtained the features related to the region of interest from the occluded portion through cross-attention.
+Park et al.
+park2022handoccnet made full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
 2.3 State Space Model
 With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences.
-In the field of natural language processing, Albert Gu et al.
-gu2022efficiently proposed the S4 model to optimize the state space model (SSM) by simulation in order to solve the problems of computational efficiency and long sequence dependency.
-However, the S4 model would process all inputs in exactly the same way, which is not in line with the practical intuition that each sequence should receive different attention.
-Therefore, Albert Gu et al.
-gu2023mamba introduced a selection mechanism to propose the Mamba architecture (S6) to learn the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences.
-Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to linear.
-Influenced by this idea, Yue Liu et al.
-ma2024umamba introduced it to vision by combining the VIT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity.
-This approach achieves linear complexity without sacrificing the global receptive field, as shown in Figure.1(c).
-In addition, some researchers ruan2024vmunet; hu2019squeezeandexcitation have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information.
-2.2 Additional information based Methods
-It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations.
-In recent years, some scholars Wang_2023; jiang2023a2jtransformer combined human a priori knowledge to guide the network to focus on the target region, Wang et al.
-Wang_2023 proposed HandGCAT, which learns the a priori knowledge of the hand joint positions through the Knowledge-Guided Graph Convolution module, and then through the Cross-Attention Transformer to guide the network to inject joint information into the occluded region.
-In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose, therefore, Qi et al.
-qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
-2.3 Attention based Methods
-Attentional mechanisms have achieved great success in the field of deep learning, thanks to their ability to map long-term contextual relevance, and in the task of hand pose estimation, occluded hand features can be learned through global mapping, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing.
-In establishing the global context of hand-object interaction, Pavlakos et al.
-pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the VIT architecture.Cho et al.
-cho2022crossattention obtain the features related to the region of interest from the occluded portion through cross-attention.
-Park et al.
-park2022handoccnet make full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
-Figure 2: Overall architecture of HandSSCA including Trunk, SSCA and Regressor.Our HandSSCA uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer.
-It then uses SSCA to obtain hand region features by pairwise spatial and channel scanning.
+In the field of natural language processing, Gu et al.
+gu2022efficiently proposed the S4 model to optimize the state space model (SSM) in order to solve the problems of computational efficiency and long sequence dependency.
+However, the S4 model would process all inputs in the exactly same way, which is not in line with the practical intuition that each sequence should receive different attention.
+Therefore, Gu et al.
+gu2023mamba introduced a selection mechanism to the Mamba architecture (S6) for learning the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences.
+Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to a linear one.
+Influenced by this idea, Liu et al.
+liu2024vmamba introduced it to vision tasks by combining the ViT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity.
+This is made possible by the selectivity of the state space model, which allows it to enhance similar features in the hidden space.
+Therefore, we abstract the selective state-space model and the cross-scanning module therein into a state space attention (SSA) module, as shown in Figure.
+1 (c).
+In addition, some researchers ruan2024vmunet; ma2024umamba; Deng2024CUMambaSS have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information.
+In this paper, we introduce state space model into the field of hand mesh reconstruction for the first time , and design a state space spatial channel attention module (S3C).
+The module performs parallel scanning in both spatial and channel dimensions, and uses a selective state-space model to achieve feature selection in the hidden space.
+This method is able to extend the effective receptive field with a small number of parameters and focuses the network attention on the hand region, thus improving the accuracy of hand mesh reconstruction.
+Figure 2: Overall architecture of HandS3C including Backbone, S3C and Regressor.
+HandS3C uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer.
+It then uses S3C to obtain hand region features through parallel spatial and channel scanning.
 Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters.
 The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
-3 Method
+3 Methodology
 3.1 Overview
-The overall network architecture is shown in Figure.2 and consists of Backbone, SSCA and Regressor components.
+The overall network architecture is shown in Fig.
+2 and consists of Backbone, S3C and Regressor components.
 Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$.
-First, the image information is extracted by ResNet-50 and the multi-scale features are fused layer by layer to get the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$.
-Then, it is fed into the SSCA module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain hand features $H\in\mathbb{R}^{256\times 32\times 32}$.
-Finally, the hand features $H$ are passed to Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates and output the MANO parameters through the FC layer to obtain the final 3D hand mesh.
+First, the appearance information is extracted by ResNet-50 based encoder-decoder and the multi-scale features are fused layer by layer to get the original feature map $F\in\mathbb{R}^{256\times 32\times 32}$.
+Then, it is fed into the S3C module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain enhanced hand features $H\in\mathbb{R}^{256\times 32\times 32}$.
+Finally, the hand features $H$ are passed to the Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates.
+For the final 3D hand mesh, we used the most popular MANO parametric hand MANO:SIGGRAPHASIA:2017.
+It is an articulated hand network generated from differentiable functions that takes as input the parameters that control the shape $\beta\in\mathbb{R}^{10}$ and pose $\theta\in\mathbb{R}^{48}$ of the hand.
 3.2 Backbone
-In the feature extraction part we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature.
+In the original feature extraction part, we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature in an U-shape encoder-decoder structure.
 Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$.
 In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling.
-Then, up-sampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of features of size $64\times 64\times 256$.
-Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the image information through average pooling.
-3.3 State Space Channel Attention
+Then, upsampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of feature extraction with size of $256\times 64\times 64$.
+Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the feature information through average pooling.
+3.3 State Space Spatial Channel Attention Module
 State Space Models.
@@
 The S6 module proposed by Gu et al.
-gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selectable scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity.
-The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{C}^{N\times N},B,C\in\mathbb{C}^{N},D\in\mathbb{C}^{1}$ with the following equation:
-$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\\ \mathrm{y(t)=Ch(t)+Dx(t)}$ (1)
-In order to achieve computationally efficient operations zimmermann2019freihand, the above ordinary differential equations need to be transformed into discrete functions.
+gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selective scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity.
+The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{R}^{N\times N},B,C\in\mathbb{R}^{N},D\in\mathbb{R}^{d}$ with the following equation:
+$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\;y(t)=Ch(t)+Dx(t)$ (1)
+In order to computationally achieve efficient operations gu2021combining, the ordinary differential equations above need to be transformed into discrete functions.
 Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times d}}$, the discretization is performed following the following rule:
 $h_{i}=\overline{A}h_{i-1}+\overline{B}x_{i},\\ y_{i}=\overline{C}h_{i}+\overline{D}x_{i},\\ \overline{A}=e^{\Delta A},\\ \overline{B}=(e^{\Delta A}-I)A^{-1}B,\\ \overline{C}=C\$ (2)
-where $\mathrm{B,C\in\mathbb{C}^{D\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{D}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series ref16:
+where $\mathrm{B,C\in\mathbb{R}^{d\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{d}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series gu2023mamba:
 $B=(e^{\Delta A}-I)A^{-1}B\approx(\Delta A)(\Delta A)^{-1}\Delta B=\Delta B$ (3)
-Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B\in\mathbb{R}^{B\times L\times N}\quad,C\in\mathbb{R}^{B\times L\times N}$, and $\Delta\in\mathbb{R}^{B\times L\times D}$ can be obtained from inputs of size $X\in\mathbb{R}^{B\times L\times D}$.
+Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B,C\in\mathbb{R}^{bs\times L\times N}$, and $\Delta\in\mathbb{R}^{bs\times L\times d}$ can be obtained from inputs of size $X\in\mathbb{R}^{bs\times L\times d}$.
 This means that S6 can sense contextual information from the inputs and update the weights.
-Space select.
+Spatial-wise Selective Scanning.
 Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module.
@@
 This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane.
-Then, the sequences from each of the four directions are passed through the S6 block to compute the feature map.
-Finally, the four sequences are merged to obtain a new sequence of the same size as the input.
-Channel select.
+Sequences in each of the four directions are computed in the Spatial-wise Selective Scanning to obtain $B,C\in\mathbb{R}^{bs\times N\times L_{H\times W}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{H\times W}}$, which in turn computes the feature map.
+Finally, the four sequences are merged to obtain a new sequence with the same size as the input sequence.
+Channel-wise Selective Scanning.
 Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba.
-However, in deep image features, channels contain richer and diverse contextual information, which is mostly ignored by these methods.
-Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Figure.3.
-We perform scan expansion and scan fusion on the $H\times C$ plane.
-First, the input features are expanded on the $H\times C$ plane and scanned from four directions to obtain the sequence $C_{1},C_{2},C_{3},C_{4}$.
-Then, they are fed into the S6 block to compute the contextual correlation in the sequence by Eq.
-(3), respectively.
-Finally, the sequences are merged and reconstructed into image feature maps of the same size.
+However, in deep features, channels contain richer and diverse contextual information, which is mostly ignored by these methods.
+Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Fig.
+3.
+We perform scan expansion and scan fusion on the $ch\times H$ plane.
+First, the input features are expanded on the $ch\times H$ plane and scanned from four directions to obtain the sequence $X_{1},X_{2},X_{3},X_{4}$.
+Then, in Channel-wise Selective Scanning, $B,C\in\mathbb{R}^{bs\times N\times L_{ch\times H}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{ch\times H}}$ is calculated by Eq.
+1-3 as an indication of the correlation between the feature channels.
+Finally, the sequences are merged and reconstructed into feature maps with the same size for images.
 Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information.
 In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field.
-We visualized the range of effective receptive fields for different attentions, as shown in Figure.1.
-As can be seen in Figure.1(d), this novel state space channel attention is able to maintain a larger and deeper receptive field range and global feature mapping capability with a reduced number of parameters and linear complexity.
+We visualized the range of effective receptive fields for different attentions, as shown in Fig.
+1.
+As can be seen in Fig.
+1 (d), the novel state space spatial channel attention module is able to maintain a larger and deeper receptive field and global feature mapping capability with a reduced number of parameters and linear complexity.
 This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
-Figure 3: SSCA module framework, we send the image features to parallel spatial and channel branches for scanning, and then process the spatial and channel resultant features and merge them through S6 block respectively, and finally sum them to get the 2D feature map.
+Figure 3: S3C module.
+In the S3C module, image features are fed into Spatial-wise Selective Scanning and Channel-wise Selective Scanning respectively.
+The correlation of spatial and channel dimensions in the hidden space is computed by selective state space modeling respectively.
+Eventually, they are merged to activate hand region features to obtain a feature map.
 3.4 Regressor
-In Regressor, we input the hand attention features obtained from SSMA.
+In Regressor, we input the hand attention features obtained from S3C into it.
 First, a 2D heat map of each joint $H$ is output through a simple Conv cascade.
-Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks [34].
+Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks he2015deep.
 Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO.
 At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
-To train our proposed HandSSCA hand pose estimation model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
+To train our proposed HandS3C hand mesh reconstruction model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
 $\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|+\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$ (4)
@@
 Junhyeong et al. cho2022crossattention 153.0M 6.5 0.982
-HandSSCA(Ours) 31.8M 7.6 7.0 0.690 0.982
+HandS3C (Ours) 31.8M 7.6 7.0 0.690 0.982
 Table 2: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
@@
 HandOccNet park2022handoccnet 39.3M 14.04 5.80
+H2ONet 10204388 - 14.00 5.70
+Feng et al. Feng20243DHR - 13.20 5.60
 HandGCAT Wang_2023 75.8M 13.76 5.60
-HandSSCA(Ours) 31.8M 12.81 5.52
+HandS3C (Ours) 31.8M 12.81 5.52
 Table 3: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
@@
 Keypoint Hampali2021KeypointTS 56.5M 10.8 0.786
-HandSSCA(Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.955
+HandS3C (Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.955
 4 Experiments
 4.1 Implementation Details
-All implementations were done in PyTorch.
-We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 learning rate to train with a batch size of 32.
-Both our training and testing were performed on a server with a CPU15 vCPU AMD and an RTX A5000 (24GB).
+All implementations were performed with PyTorch.
+We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 to train the model with batch size of 32.
+Both our training and testing were performed on a server with an GPU RTX A5000 (24GB).
 All other details will be provided in our code.
+Figure 4: Qualitative results for FREIHAND and HO3D.
+From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, and the hand mesh.
 4.2 Datasets and Evaluation Metrics
@@
 For FreiHand, we report PA-MPJPE and PA-MPVPE in $mm$ as well as F-scores.
-For HO3D we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server.
-For DexYCB we report MPJPE and PA-MPJPE in mm.
+For HO3D, we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server.
+For DexYCB, we report MPJPE and PA-MPJPE in $mm$.
 Table 4: ABLATION RESULTS FOR THE FREIHAND.
@@
 Baseline+Self-Attention 7.7 7.0 0.688 0.981
-Baseline+SSCA 7.6 7.7 0.690 0.982
+Baseline+S3C 7.6 7.0 0.690 0.982
 Table 5: ABLATION RESULTS FOR THE DEXYCB.
@@
 Baseline+Self-Attention 14.40 5.81
-Baseline+SSCA 12.81 5.52
+Baseline+S3C 12.81 5.52
 Table 6: ABLATION RESULTS FOR THE HO3D.
@@
 Baseline+SSA (0.57M) 9.8 9.8 0.521 0.952
-Baseline+SSCA(0.58M) 9.5 9.5 0.518 0.955
+Baseline+serial-S3C (0.58M) 9.9 9.8 0.516 0.952
+Baseline+S3C (0.58M) 9.5 9.5 0.518 0.955
 4.3 Comparisons with the State-of-the-art Methods
-We compare our proposed method with several recent state-of-the-art methods on three datasets, FreiHand, DEXYCB, and HO3D, as shown in Tables 1-3.
+We compare our proposed method with several recent state-of-the-art methods on three datasets as shown in Tables 1-3.
 As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830.
@@
 liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced.
-To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023.
+To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023; 10204388; Feng20243DHR.
 In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included.
-Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Figure.4.
+Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Fig.4 - 5.
+Figure 5: Qualitative results for DEXYCB.
+From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, the hand mesh and HandOCC park2022handoccnet hand mesh results.
+Figure 6: Attention characterization map.
+It is worth noting that the S3C module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region.
+And the method expands the range of the receptive field so that it can capture the global feature mapping.
+(Orange is the occluded region and red is the hand interaction edge region)
 4.4 Ablation Study
-In order to validate the effectiveness of the HandSSCA method, we further analyzed the different attention modules.
+In order to validate the effectiveness of the HandS3C method, we further analyze the different attention modules.
 The results of the ablation experiments on different datasets are shown in Tables 4-6.
-On the FreiHand and DEXYCB datasets, we designed three experiments, Baseline, Adding Self-Attention Module, and Adding SSCA Module, as a way to validate that our proposed SSCA is able to expand the effective receptive field while enhancing the hand features.
-On the HO3D dataset, we designed four experiments, which are Baseline, adding the self-attention module, adding the traditional state space attention module (SSA) and adding the state space channel attention module (SSCA).
-In this way, we demonstrate that the state space model with the addition of channel information can enhance the ability to extract hand features in the occluded state while having a wider effective receptive field.
+On the FreiHand and DEXYCB datasets, we perform three experiments, i.e., Baseline, Adding Self-Attention Module, and Adding S3C Module, as a way to validate that our proposed S3C is able to expand the effective receptive field while enhancing the hand features.
+On the HO3D dataset, we conducted five experiments: baseline, adding a self-attention module, adding a traditional state-space attention module (SSA), adding a serial state-space spatial-channel attention module (serial-S3C), and adding a parallel state-space spatial-channel attention module (S3C).
+With this approach, we demonstrate that a state-space model with added channel information enhances the ability to extract hand features under occlusion, with a wider effective receptive field, and the effectiveness of the parallel structure.
 From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information.
-In Figure.5, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges.
+In Fig.
+6, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges.
 The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent.
-The SSCA module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes.
+The S3C module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes.
 Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
@@
 However, observing other metrics will show that the channel dimension in the image will then contain richer feature information, which happens to be ignored by SSA.
-Figure 4: Qualitative results for the three datasets.
-From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by SSCA, the heat map of keypoints in Regressor, and the hand mesh.
-Figure 5: Attention characterization map.
-It is worth noting that the SSCA module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region.
-And the method expands the range of the receptive field so that it can capture the global feature mapping.
-(Orange is the occluded region and red is the hand interaction edge region)
+For the choice of serial and parallel structures, success was achieved in image restoration through serial structures in the study of Deng et al.
+Deng2024CUMambaSS.
+However, since SSM resembles the linear computation of recurrent neural networks (RNN).
+Serial scanning in space and channel add extra computation time.
+Meanwhile, compared to the parallel structure, the serial structure does not balance the contributions of spatial and channel dimensions well, as shown in the ablation experiments in Table 6.
 5 Conclusion
-In order to effectively mitigate the challenges of the hand pose estimation task under occlusion conditions, we introduce state-space modeling into the field of hand pose estimation for the first time and propose the HandSSCA method.
-The method expands the effective receptive field and acquires hand features under occlusion by parallel cross-scanning in the spatial plane and the channel plane.
+In order to effectively mitigate the challenges of the hand mesh reconstruction task under occlusion conditions, we introduced the state space model into the task of hand mesh reconstruction for the first time and propose the HandS3C method.
+The method expanded the effective receptive field and acquired hand features under occlusion by parallel cross-scanning in the spatial and channel planes.
 Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection.
@@
 Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations.
-Meanwhile, we expect this spatial and channel-parallel state-selective attention architecture to make progress in different downstream tasks.
+Meanwhile, we expect this spatial and channel-wise state-space attention architecture to make progress in different downstream tasks.
+[…]

### v4

@@
 Most previous works attempt to explore more additional information and adopt attention mechanisms for improving 3D reconstruction performance, while it would increase computational complexity simultaneously.
-To achieve a performance-reserving architecture with high computational efficiency, in this work, we propose a simple but effective 3D hand mesh reconstruction network (i.e., HandS3C), which is the first time to incorporate state space model into the task of hand mesh reconstruction.
+To achieve a performance preserving architecture with high computational efficiency, in this work, we propose a simple but effective 3D hand mesh reconstruction network (i.e., HandS3C), which is the first time to incorporate state space model into the task of hand mesh reconstruction.
 In the network, we design a novel state-space spatial-channel attention module that extends the effective receptive field, extracts hand features in the spatial dimension, and enhances regional features of hands in the channel dimension.
 This helps to reconstruct a complete and detailed hand mesh.
-Extensive experiments conducted on well-known datasets facing heavy occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandS3C achieves state-of-the-art performance while maintaining a minimal parameters.
-1 Introduction
-The reconstruction of 3D hand mesh from one single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction.
-3D hand mesh reconstruction based on RGB images for the reconstruction has attracted much attention due to its low cost and computational friendliness.
-Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field.
+Extensive experiments conducted on well-known datasets facing heavy occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandS3C achieves state-of-the-art performance while maintaining minimal parameters.
+Code can be available in https://github.com/JiaoZixun/HandS3C.
+Index Terms:
+3D Hand Mesh Reconstruction, Deep Learning, Effective Receptive Field, Human-computer Interaction, State Space Model.
+I Introduction
+The reconstruction of 3D hand mesh from one single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction and has attracted much attention due to its low cost and computational friendliness.
+Influenced by the rapid development of deep learning techniques, many excellent works [1, 2, 3, 4, 7, 5, 6] have emerged for this task.
 These works focus on important issues related to hand-object occlusion and the loss of local information.
-Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when the prior knowledge is not sufficient.
+Despite significant advancements, most methods still struggle to reconstruct hand poses in some scenarios with substantial occlusion or when the prior knowledge is not sufficient.
 The previous methods can be broadly categorized into two categories.
-One is to incorporate additional prior knowledge.
-To cite few, Zimm et al.
-zimmermann2017learning obtained hand region features via a hand segmentation network.
-Wang et al.
-Wang_2023 added keypoint coordinates as strong prior information to constrain hand pose.
-These methods can enhance pose reconstruction performance, but lack essential prior information in practical applications.
-Another category is to learn the intrinsic connection between image features through deep networks, Park et al.
-park2022handoccnet extracted the complete hand information by distinguishing the features.
-Jiao et al.
-https://doi.org/10.1049/ell2.13040 used the Transformer architecture to obtain the global information and expect to learn the hand feature mapping from the occluded part.
-These methods tend to mine the association between image features, but there are still some limitations.
+One is to incorporate additional prior knowledge [7, 8, 9, 10].
+To cite few, Wang et al.
+[7] added keypoint coordinates as strong prior information to constrain hand pose.
+Zimmermann et al.
+[8] obtained hand region features via a hand segmentation network.
+Xie et al.
+[9] combined physiological aspects of hand motion to generate kinematic-compliant hand postures.
+These methods can enhance pose reconstruction performance, but lack essential prior information in practical applications (e.g., VR/AR games, robotics).
+Another class of approaches is to learn potential associations of image features through deeper architectures, but inevitably there is a scarcity of real annotation data.
+Some researchers [5, 11, 12] use Transformer architecture to obtain global information and expect to learn hand feature mapping from occluded parts.
+Other researchers [13] aggregated a large number of real datasets containing annotations for exploring associations between image features.
+However, these studies still have some limitations.
 First, the attention mechanism leads to excessive computational complexity.
-Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure.
-1 (b).
+Second, an overload of global information can interfere with the effective receptive field [14].
+Figure 1: Overall architecture of HandS3C including Backbone, S3C and Regressor.
+In Backbone, ResNet-50 based encoder-decoder is employed to extract image features and fuses multiscale features.
+In S3C, image features are fed into Spatial-wise Selective Scanning and Channel-wise Selective Scanning modules, respectively.
+In Regressor, heat map features are obtained using convolution operation .
+Further MANO parameters are obtained from the FC layer to generate the hand mesh.
 Based on the above observations, we propose a simple but effective 3D hand mesh reconstruction network called as HandS3C.
-This framework incorporates the state space model liu2024vmamba; ma2024umamba; ruan2024vmunet into the field of hand mesh reconstruction for the first time.
-At the same time, combining the idea of spatial-channel feature fusion hu2019squeezeandexcitation, State Space Spatial Channel (S3C) attention module is constructed to emphasize the contribution of channel-dimensional features.
+This framework incorporates the state space model [15, 16, 17, 18] into the field of hand mesh reconstruction for the first time.
+At the same time, combining the idea of spatial-channel feature fusion [19], State Space Spatial Channel (S3C) attention module is constructed to emphasize the contribution of channel-dimensional features.
 This not only increases the range of effective receptive fields in the network, but also effectively avoids the square-step computational complexity of traditional attention.
 Specifically, our main contributions are summarized as follows:
-(1) We design a 3D hand mesh reconstruction network called as HandS3C, which introduces the state space model to the task of hand mesh reconstruction for the first time.
+  • We design a 3D hand mesh reconstruction network called as HandS3C, which introduces the state space model to the task of hand mesh reconstruction for the first time.
 This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
-(2) We propose a spatial and channel-wise parallel scanning approach to compensate for the missing channel information in planar scanning.
-The state-space spatial-channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
-(3) Our method achieves state-of-the-art on three datasets, FREIHAND, DEXYCB and HO3D, with a small number of parameters.
-Figure 1: Effective receptive fields (from left to right: FPN, Self-Attention, State Space Attention (SSA) and State-Space Spatial-Channel Attention (S3C)).
-2 Related work
-2.1 Additional Information based Methods
-It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations.
-In recent years, some scholars Wang_2023; jiang2023a2jtransformer incorporate human prior knowledge to guide the network to focus on the target region, Wang et al.
-Wang_2023 proposed HandGCAT, which learns the prior knowledge of the hand joint positions through the knowledge-guided graph convolution module, and then through the cross-attention Transformer to guide the network to inject joint information into the occluded region.
-In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose.
-To end this, Qi et al.
-qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
-2.2 Attention based Methods
-Attention mechanisms have achieved great success in the field of deep learning, due to their ability to map long-term contextual relevance.
-In the task of hand mesh reconstruction, occluded hand features can be learned through global mapping with various attention modules, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing.
-In establishing the global context of hand-object interaction, Pavlakos et al.
-pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the Vision Transformer (ViT) architecture.
-Cho et al.
-cho2022crossattention obtained the features related to the region of interest from the occluded portion through cross-attention.
-Park et al.
-park2022handoccnet made full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
-2.3 State Space Model
-With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences.
-In the field of natural language processing, Gu et al.
-gu2022efficiently proposed the S4 model to optimize the state space model (SSM) in order to solve the problems of computational efficiency and long sequence dependency.
-However, the S4 model would process all inputs in the exactly same way, which is not in line with the practical intuition that each sequence should receive different attention.
-Therefore, Gu et al.
-gu2023mamba introduced a selection mechanism to the Mamba architecture (S6) for learning the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences.
-Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to a linear one.
-Influenced by this idea, Liu et al.
-liu2024vmamba introduced it to vision tasks by combining the ViT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity.
-This is made possible by the selectivity of the state space model, which allows it to enhance similar features in the hidden space.
-Therefore, we abstract the selective state-space model and the cross-scanning module therein into a state space attention (SSA) module, as shown in Figure.
-1 (c).
-In addition, some researchers ruan2024vmunet; ma2024umamba; Deng2024CUMambaSS have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information.
-In this paper, we introduce state space model into the field of hand mesh reconstruction for the first time , and design a state space spatial channel attention module (S3C).
-The module performs parallel scanning in both spatial and channel dimensions, and uses a selective state-space model to achieve feature selection in the hidden space.
-This method is able to extend the effective receptive field with a small number of parameters and focuses the network attention on the hand region, thus improving the accuracy of hand mesh reconstruction.
-Figure 2: Overall architecture of HandS3C including Backbone, S3C and Regressor.
-HandS3C uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer.
-It then uses S3C to obtain hand region features through parallel spatial and channel scanning.
-Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters.
-The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
-3 Methodology
-3.1 Overview
+  • We propose a spatial and channel-wise parallel scanning approach to compensate for the missing channel information in planar scanning.
+The state-space spatial-channel attention module is constructed, which can enhance the effective receptive field range while maintaining small number of parameters.
+  • Our method achieves state-of-the-art performance on three datasets, FREIHAND, DEXYCB and HO3D, with smaller number of parameters.
+II Methodology
+II-A Overview
 The overall network architecture is shown in Fig.
 1 and consists of Backbone, S3C and Regressor components.
-Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$.
-First, the appearance information is extracted by ResNet-50 based encoder-decoder and the multi-scale features are fused layer by layer to get the original feature map $F\in\mathbb{R}^{256\times 32\times 32}$.
+Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$, the appearance information is firstly extracted by ResNet-50 based encoder-decoder and the multi-scale features are fused layer by layer to get the original feature map $F\in\mathbb{R}^{256\times 32\times 32}$.
 Then, it is fed into the S3C module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain enhanced hand features $H\in\mathbb{R}^{256\times 32\times 32}$.
 Finally, the hand features $H$ are passed to the Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates.
-For the final 3D hand mesh, we used the most popular MANO parametric hand MANO:SIGGRAPHASIA:2017.
+For the final 3D hand mesh, we use the most popular MANO parametric hand [20].
 It is an articulated hand network generated from differentiable functions that takes as input the parameters that control the shape $\beta\in\mathbb{R}^{10}$ and pose $\theta\in\mathbb{R}^{48}$ of the hand.
-3.2 Backbone
-In the original feature extraction part, we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature in an U-shape encoder-decoder structure.
+II-B Backbone
+In the original feature extraction part, we use ResNet-50 to extract image information and fuse multi-scale features layer by layer in an U-shape encoder-decoder structure.
 Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$.
@@
 Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the feature information through average pooling.
-3.3 State Space Spatial Channel Attention Module
-State Space Models.
-S4ND nguyen2022s4nd was the first study to introduce state space models into the field of vision by enabling it to receive 2D image information through a simple extension of the S4 model.
-The S6 module proposed by Gu et al.
-gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selective scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity.
-The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{R}^{N\times N},B,C\in\mathbb{R}^{N},D\in\mathbb{R}^{d}$ with the following equation:
-$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\;y(t)=Ch(t)+Dx(t)$ (1)
-In order to computationally achieve efficient operations gu2021combining, the ordinary differential equations above need to be transformed into discrete functions.
-Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times d}}$, the discretization is performed following the following rule:
-$h_{i}=\overline{A}h_{i-1}+\overline{B}x_{i},\\ y_{i}=\overline{C}h_{i}+\overline{D}x_{i},\\ \overline{A}=e^{\Delta A},\\ \overline{B}=(e^{\Delta A}-I)A^{-1}B,\\ \overline{C}=C\$ (2)
-where $\mathrm{B,C\in\mathbb{R}^{d\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{d}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series gu2023mamba:
-$B=(e^{\Delta A}-I)A^{-1}B\approx(\Delta A)(\Delta A)^{-1}\Delta B=\Delta B$ (3)
-Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B,C\in\mathbb{R}^{bs\times L\times N}$, and $\Delta\in\mathbb{R}^{bs\times L\times d}$ can be obtained from inputs of size $X\in\mathbb{R}^{bs\times L\times d}$.
-This means that S6 can sense contextual information from the inputs and update the weights.
-Spatial-wise Selective Scanning.
+II-C State Space Spatial Channel (S3C) Attention Module
+Spatial-wise Selective Scanning
 Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module.
-The 1-D convolution is extended to 2-D convolution in S4ND gu2021combining, and this approach leads to the loss of the ability to dynamically update weights based on context.
-In VMamba liu2024vmamba, rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module.
+The 1-D convolution is extended to 2-D convolution in S4ND [21], and this approach leads to the loss of the ability to dynamically update weights based on context.
+In VMamba [15], rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module.
 This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane.
@@
 Finally, the four sequences are merged to obtain a new sequence with the same size as the input sequence.
-Channel-wise Selective Scanning.
-Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba.
+Channel-wise Selective Scanning
+Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al [15].
 However, in deep features, channels contain richer and diverse contextual information, which is mostly ignored by these methods.
 Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Fig.
-3.
+1.
 We perform scan expansion and scan fusion on the $ch\times H$ plane.
@@
 Finally, the sequences are merged and reconstructed into feature maps with the same size for images.
-Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information.
+Image features are spatially selected by activating the mapping of each pixel in the four directions to obtain local texture and structure information.
 In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field.
-We visualized the range of effective receptive fields for different attentions, as shown in Fig.
-1.
-As can be seen in Fig.
-1 (d), the novel state space spatial channel attention module is able to maintain a larger and deeper receptive field and global feature mapping capability with a reduced number of parameters and linear complexity.
+The novel state space spatial channel attention module is able to maintain a larger and deeper receptive field and global feature mapping capability with a reduced number of parameters and linear complexity.
 This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
-Figure 3: S3C module.
-In the S3C module, image features are fed into Spatial-wise Selective Scanning and Channel-wise Selective Scanning respectively.
-The correlation of spatial and channel dimensions in the hidden space is computed by selective state space modeling respectively.
-Eventually, they are merged to activate hand region features to obtain a feature map.
-3.4 Regressor
+$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\;y(t)=Ch(t)+Dx(t)$ (1)
+$h_{i}=$ $\overline{A}h_{i-1}+\overline{B}x_{i},y_{i}=Ch_{i}+Dx_{i},$ (2)
+$\overline{A}=$ $e^{\Delta A},\overline{B}=(e^{\Delta A}-I)A^{-1}B,\overline{C}=C$
+$B=(e^{\Delta A}-I)A^{-1}B\approx(\Delta A)(\Delta A)^{-1}\Delta B=\Delta B$ (3)
+II-D Regressor
 In Regressor, we input the hand attention features obtained from S3C into it.
-First, a 2D heat map of each joint $H$ is output through a simple Conv cascade.
-Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks he2015deep.
-Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO.
-At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
-To train our proposed HandS3C hand mesh reconstruction model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
-$\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|+\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$ (4)
+First, the 2D joint heat map $HM$ is computed using a convolution operation.
+Then, a cascade of hand attention features and a 2D heatmap $HM$ is input into four residual blocks.
+Finally, the pose parameter $\theta\in\mathbb{R}^{48}$ and the shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO are predicted.
+Detailed 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ are obtained from the MANO hand model generated.
+To train our proposed HandS3C model for hand mesh reconstruction, we constrain the model by calculating L2 loss.
+$\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-$ $\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|$ (4)
+$+$ $\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$
 where $*^{\prime}$ denotes the true value, * denotes the predicted value, and $\alpha_{\theta}=10,\alpha_{\beta}=0.1,\alpha_{2d}=1\mathrm{e}2,\alpha_{3d}=1\mathrm{e}4,\alpha_{V}=1\mathrm{e}4$.
 the optimal solution is obtained by minimizing this objective function for training.
-Table 1: COMPARISON WITH STATE-OF-THE-ART METHODS ON FREIHAND.
-Method Params↓ PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
-Hasson et al. Hasson2019LearningJR 33.5M 13.3 13.3 0.429 0.907
-ExPose Choutas2020MonocularEB 11.8 12.2 0.484 0.918
-Kulon et al. kulon2020weaklysupervised 519.0M 8.4 8.6 0.614 0.966
-I2L-MeshNet moon2020i2lmeshnet 136.8M 7.4 7.6 0.681 0.973
-Pose2Mesh choi2021pose2mesh 75.0M 7.4 7.6 0.683 0.973
-Tang et al. 9710830 156.5M 6.7 6.7 0.724 0.981
-Junhyeong et al. cho2022crossattention 153.0M 6.5 0.982
-HandS3C (Ours) 31.8M 7.6 7.0 0.690 0.982
-Table 2: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
+TABLE I: COMPARISON WITH STATE-OF-THE-ART METHODS ON FREIHAND.
+Method Params↓ PA- PA- F@5↑ F@15↑
+MPJPE↓ MPVPE↓
+Hasson et al. [1] 33.5M 13.3 13.3 0.429 0.907
+ExPose [3] 11.8 12.2 0.484 0.918
+Kulon et al. [22] 519.0M 8.4 8.6 0.614 0.966
+I2L-MeshNet [23] 136.8M 7.4 7.6 0.681 0.973
+Pose2Mesh [24] 75.0M 7.4 7.6 0.683 0.973
+Tang et al. [25] 156.5M 6.7 6.7 0.724 0.981
+Junhyeong et al. [26] 153.0M 6.5 0.982
+AHRNET [27] 7.2 0.978
+Baseline 8.0 7.4 0.669 0.977
+Baseline+Self-Attention 7.7 7.0 0.688 0.981
+HandS3C (Ours) 31.8M 7.0 6.3 0.727 0.989
+TABLE II: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
 Method Params↓ MPJPE↓ PA-MPJPE↓
-Spurr et al. spurr2020weakly 17.34 6.83
-METRO lin2021endtoend 229.5M 15.24 6.99
-Liu et al. liu2021semisupervised 34.9M 15.28 6.58
-HandOccNet park2022handoccnet 39.3M 14.04 5.80
-H2ONet 10204388 - 14.00 5.70
-Feng et al. Feng20243DHR - 13.20 5.60
-HandGCAT Wang_2023 75.8M 13.76 5.60
+Spurr et al. [28] 17.34 6.83
+METRO [29] 229.5M 15.24 6.99
+Liu et al. [30] 34.9M 15.28 6.58
+HandOccNet [5] 39.3M 14.04 5.80
+H2ONet [31] - 14.00 5.70
+Feng et al. [32] - 13.20 5.60
+HandGCAT [7] 75.8M 13.76 5.60
+MS-MANO [9] 12.92
+Baseline 14.07 5.78
+Baseline+Self-Attention 14.40 5.81
 HandS3C (Ours) 31.8M 12.81 5.52
-Table 3: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
+TABLE III: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
 Method Params↓ PA-MPJPE↓ PA-MPJPE AUC↑ PA-MPVPE↓ PA-MPVPE AUC↑ F@5↑ F@15↑
-I2L-MeshNet moon2020i2lmeshnet 136.8M 11.2 0.775 13.9 0.722 0.409 0.932
-Hasson et al. hasson2020leveraging 13.2M 11.0 0.780 11.2 0.777 0.464 0.939
-Hampali et al. hampali2020honnotate 10.7 0.788 10.6 0.790 0.506 0.942
-METRO lin2021endtoend 229.5M 10.4 0.792 11.1 0.779 0.484 0.946
-Liu et al. liu2021semisupervised 34.9M 9.9 0.803 9.5 0.810 0.528 0.956
-I2UV-HandNet chen2021i2uvhandnet 9.9 0.804 10.1 0.799 0.500 0.943
-ArtiBoost li2022artiboost 25.3M 11.4 0.773 10.9 0.782 0.488 0.944
-Keypoint Hampali2021KeypointTS 56.5M 10.8 0.786
-HandS3C (Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.955
-4 Experiments
-4.1 Implementation Details
-All implementations were performed with PyTorch.
-We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 to train the model with batch size of 32.
-Both our training and testing were performed on a server with an GPU RTX A5000 (24GB).
+I2L-MeshNet [23] 136.8M 11.2 0.775 13.9 0.722 0.409 0.932
+Hasson et al. [38] 13.2M 11.0 0.780 11.2 0.777 0.464 0.939
+Hampali et al. [35] 10.7 0.788 10.6 0.790 0.506 0.942
+METRO [29] 229.5M 10.4 0.792 11.1 0.779 0.484 0.946
+Liu et al. [30] 34.9M 9.9 0.803 9.5 0.810 0.528 0.956
+I2UV-HandNet [36] 9.9 0.804 10.1 0.799 0.500 0.943
+ArtiBoost [37] 25.3M 11.4 0.773 10.9 0.782 0.488 0.944
+Keypoint [4] 56.5M 10.8 0.786
+Baseline 9.8 0.798 9.7 0.801 0.516 0.949
+Baseline+Self-Attention(4.6M) 10.1 0.798 10.1 0.798 0.502 0.947
+Baseline+SSA(0.57M) 9.8 0.803 9.8 0.804 0.521 0.952
+Baseline+serial-S3C(0.58M) 9.9 9.8 0.516 0.952
+Baseline+S3C(0.58M) (Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.957
+III Experiments
+III-A Implementation Details
+We train on three publicly available datasets, FREIHAND[33], DEXYCB[34], and HO3D[35], and the model was trained by annealing every 10th time starting from an initial learning rate of 10-4 to train the model with batch size of 32.
+Our training and testing were performed on a server equipped with a GPU RTX A5000 (24GB).
 All other details will be provided in our code.
+In the FREIHAND, we report PA-MPJPE and PA-MPVPE and F-scores in $mm$.
+For DEXYCB, we report MPJPE and PA-MPJPE in $mm$.
+For HO3D, we report PA-MPJPE/MPVPE, AUC and F scores returned from the official evaluation server.
 Figure 2: Qualitative results for FREIHAND and HO3D.
 From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, and the hand mesh.
-4.2 Datasets and Evaluation Metrics
-FREIHAND.
-This dataset zimmermann2019freihand presents the first massively diverse hand dataset containing 32,560 training samples with the possibility of replacing four backgrounds.
-Also, there are 3960 test samples under occlusion and poor lighting.
-DEXYCB.
-This dataset chao2021dexycb is a large-scale dataset for object grasping and contains 582K image frames grasped on 20 YCB objects.
-We use the official “S0” segmentation to evaluate the right hand pose.
-HO3D.
-This dataset hampali2020honnotate is a hand-object interaction dataset with severe occlusion and contains 66,034 training samples and 11,524 test samples.
-The dataset provides RGB images with MANO-based hand joints and meshes and camera parameters.
-The test set provides only the annotations of the RGB images and the hand bounding box, and the test results can be evaluated through an online submission system.
-Evaluation metrics.
-For FreiHand, we report PA-MPJPE and PA-MPVPE in $mm$ as well as F-scores.
-For HO3D, we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server.
-For DexYCB, we report MPJPE and PA-MPJPE in $mm$.
-Table 4: ABLATION RESULTS FOR THE FREIHAND.
-Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
-Baseline 8.0 7.4 0.669 0.977
-Baseline+Self-Attention 7.7 7.0 0.688 0.981
-Baseline+S3C 7.6 7.0 0.690 0.982
-Table 5: ABLATION RESULTS FOR THE DEXYCB.
-Method MPJPE↓ PA-MPJPE↓
-Baseline 14.07 5.78
-Baseline+Self-Attention 14.40 5.81
-Baseline+S3C 12.81 5.52
-Table 6: ABLATION RESULTS FOR THE HO3D.
-Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
-Baseline 9.8 9.7 0.516 0.954
-Baseline+Self-Attention (4.6M) 10.1 10.1 0.502 0.947
-Baseline+SSA (0.57M) 9.8 9.8 0.521 0.952
-Baseline+serial-S3C (0.58M) 9.9 9.8 0.516 0.952
-Baseline+S3C (0.58M) 9.5 9.5 0.518 0.955
-4.3 Comparisons with the State-of-the-art Methods
-We compare our proposed method with several recent state-of-the-art methods on three datasets as shown in Tables 1-3.
-As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830.
-Moreover, its number of parameters shrinks by a factor of nearly 5 compared to models with higher accuracy cho2022crossattention.
-On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods Hampali2021KeypointTS.
+III-B Comparisons with the State-of-the-art Methods
+We compare our proposed method with several recent state-of-the-art methods on three datasets as shown in Tables I-III.
+As can be seen in table I, our method only slightly underperforms the other methods[26] in the PA-MPJPE metrics, while remaining ahead in the other metrics, and it is important to note that the amount of parameters in our method has been reduced by a factor of almost 5[26].
+On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods [4].
 Meanwhile, when the number of parameters is similar to the method of Liu et al.
-liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced.
-To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023; 10204388; Feng20243DHR.
+[30], all other metrics are greatly improved, and only the F-score metric is slightly reduced.
+To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table II, and all the metrics outperform the recent state-of-the-art models [7, 9, 31, 32].
 In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included.
-Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Fig.4 - 5.
+Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in fig.2 - 3.
 Figure 3: Qualitative results for DEXYCB.
-From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, the hand mesh and HandOCC park2022handoccnet hand mesh results.
-Figure 6: Attention characterization map.
-It is worth noting that the S3C module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region.
-And the method expands the range of the receptive field so that it can capture the global feature mapping.
-(Orange is the occluded region and red is the hand interaction edge region)
-4.4 Ablation Study
-In order to validate the effectiveness of the HandS3C method, we further analyze the different attention modules.
-The results of the ablation experiments on different datasets are shown in Tables 4-6.
-On the FreiHand and DEXYCB datasets, we perform three experiments, i.e., Baseline, Adding Self-Attention Module, and Adding S3C Module, as a way to validate that our proposed S3C is able to expand the effective receptive field while enhancing the hand features.
-On the HO3D dataset, we conducted five experiments: baseline, adding a self-attention module, adding a traditional state-space attention module (SSA), adding a serial state-space spatial-channel attention module (serial-S3C), and adding a parallel state-space spatial-channel attention module (S3C).
-With this approach, we demonstrate that a state-space model with added channel information enhances the ability to extract hand features under occlusion, with a wider effective receptive field, and the effectiveness of the parallel structure.
-From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information.
-In Fig.
-6, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges.
-The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent.
-The S3C module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes.
-Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
-In Table 6, it can be found that after the introduction of the state selection of the channel plane, the number of parameters of the model is only slightly increased by 0.01M, however, a significant improvement of nearly 2% is realized in the two performance metrics of PA-MPJPE and PA-MPVPE.
-This is due to the fact that the state selection in the channel plane enhances the selection and fusion of image features, and is able to extract more important channel information.
-The current state-space model focuses more on multi-directional scanning in the spatial plane, which is the reason for the 0.3% improvement in the F@5 metrics when using SSA.
-Using only the spatial plane for state selection, each pixel will focus more on neighboring pixels.
-However, observing other metrics will show that the channel dimension in the image will then contain richer feature information, which happens to be ignored by SSA.
-For the choice of serial and parallel structures, success was achieved in image restoration through serial structures in the study of Deng et al.
-Deng2024CUMambaSS.
+From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, the hand mesh and HandOCC [5] hand mesh results.
+III-C Ablation Study
+In order to verify the effectiveness of the HandS3C method, we further analyzed different attention modules.
+On the FreiHand and DEXYCB datasets, we conducted three experiments: baseline, adding a self-attention module, and adding an S3C module.
+On the HO3D dataset, we supplemented these three ablation experiments with experiments on the traditional state-space attention module (SSA) and the serial state-space spatial channel attention module (serial-S3C).
+With these ablation experiments, we demonstrate that the state-space model with added channel information can improve the ability to extract occluded partial hand features while extending the effective sensory field enhancement, and verify the effectiveness of the parallel structure.
+The results are summarized in Tables I-III.
+From Tables I and III, it can be seen that on smaller scale datasets, the self-attention mechanism is difficult to learn the global features of the occluded part, but instead introduces more noise information to cause interference.
+The S3C module, on the other hand, expands the range of the sensory field through the spatial plane selection mechanism, and then enhances the hand region features directionally through the selection of the channel plane, which in turn avoids the interference of global noise.
+From Table III, it can be found that compared with the original SSM, the number of parameters of the model is only slightly increased by 0.01M after the introduction of the state selection of the channel plane, but a significant enhancement of nearly 2% is realized in the two performance metrics of PA-MPJPE and PA-MPVPE.
+This is due to the fact that our approach emphasizes the importance of channel information, which in combination with spatial information can expand the effective sensory field and thus extract the features of the region of focus.
+The study of Deng et al.
+successfully realized image restoration through serial structure [39].
 However, since SSM resembles the linear computation of recurrent neural networks (RNN).
 Serial scanning in space and channel add extra computation time.
-Meanwhile, compared to the parallel structure, the serial structure does not balance the contributions of spatial and channel dimensions well, as shown in the ablation experiments in Table 6.
-5 Conclusion
-In order to effectively mitigate the challenges of the hand mesh reconstruction task under occlusion conditions, we introduced the state space model into the task of hand mesh reconstruction for the first time and propose the HandS3C method.
-The method expanded the effective receptive field and acquired hand features under occlusion by parallel cross-scanning in the spatial and channel planes.
-Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection.
-This novel attention architecture is able to acquire sufficient global information and expand the effective sensory field while keeping the number of parameters small.
-Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations.
-Meanwhile, we expect this spatial and channel-wise state-space attention architecture to make progress in different downstream tasks.
-[…]
+Meanwhile, compared to the parallel structure, the serial structure does not balance the contributions of spatial and channel dimensions well, as shown in the ablation experiments in Table III.
+IV Conclusion
+In order to effectively alleviate the challenges of the hand mesh reconstruction task under occlusion conditions, we firstly introduce the state space model into the hand mesh reconstruction task and propose the HandS3C method.
+The method enlarges the effective receptive field by performing parallel cross-scanning of the spatial plane and the channel plane.
+Thus, hand features under occlusion conditions are acquired while keeping fewer parameters.
+Extensive experimental results show that this approach can accurately predict hand pose under heavy occlusion.
+The ablation study also validates the effectiveness of the space state model in comparison to the attention-based model.
+V Acknowledgment
+This work was supported by the National Natural Science Foundation of the China, Shaanxi Provincial Key Industry Innovation Chain Program, 2024 public digital cultural service research project of the national public cultural development center of the Ministry of culture and tourism, and the scientific research program of Shaanxi Provincial Department of Education [grant numbers 62072362 and 12101479, 2020ZDLGY07-05, GGSZWHFW2024-017, 23jp060].